Is the ordered pair (-4,-9) a solution to the linear equation: y=3/4x-6?
Answer by stanbon(75887)

Is the ordered pair (-4,-9) a solution to the linear equation:
y=3/4x-6?
---
Substitute: -9 = (3/4)(-4)-6
---
-9 = -3 -6
---
-9 = -9
----
Answer: Yes
================
Cheers,
Stan H.
================